Question 2 Basic Probability: Each question is worth 2 marks - Total (10 marks) Business makes a profit Business makes a loss International Manager 0.15 0.2 Local Manager 0.35 0.3 A. P (Local Manager) B. P (Business makes a loss) C. P (International Manager OR Business makes a profit) D. P (Local Manager AND International Manager) E. P (Business makes a loss | Manager is International)
Answer #1

Ans:

profit loss Total
International 0.15 0.2 0.35
local 0.35 0.3 0.65
Total 0.5 0.5 1

a)

P(local)=0.35+0.3=0.65

b)

P(loss)=0.2+0.3=0.5

c)

P(international or profit)=P(international)+P(profit)-P(international and profit)

=0.35+0.5-0.15

=0.7

d)

P(local and international)=0

e)

P(loss/international)=0.2/0.35=0.5714